# Prompt reinforcing for long-term planning of large language models

### v1

@@ 5.2 Effect of different styles and input signals of textual-based feedbacker
(a) Different textual feedback generation methods
(b) Different info for TD-style+replay method
Figure 4: The training curves of different optimisation methods. Each setting is trained on 4 seeds and evaluated on 100 dialogues. The line is the average success and the shadow is the standard error.
The training curves of FnCTOD optimised by the methods of MC-style, TD-style, and TD-style+replay with Gemini-2.0-flash are shown in 4(a) (See results with other LLMs in Figure 6). Similar to the behaviour in traditional RL optimisation, MC-style exhibits higher variance during the early stages of training, whereas TD-style is more stable and converges faster. With further training, their final performances become comparable. In contrast, incorporating experience replay into the rewriter yields more stable training and achieves the best overall performance.
We conduct a further ablation study on the impact of different information as input to the feedbacker (as shown in 4(b)). The basic setting passes the dialogue in pure text. The subjective setting includes the user goal, and the believe setting adds the API call in comparison to the basic setting, respectively. The full setting is our proposed TD-style+replay, including both the user goal and the system API call.
@@
5.3 Prompting limitations on underrepresented topics in LLMs
In general medicine, our method consistently outperforms the fully fine-tuned HuatuoGPT-II with an 86.7% win rate and is preferred over other prompting-based baselines. On the other hand, traditional Chinese medicine is more challenging. For example, our system’s preference rate drops by 41% compared to Huatuo when transitioning from general medicine to traditional Chinese medicine. However, despite this drop in preference, our proposed method is still favoured in general.
@@ 6 Conclusions
Ethic Statement
This work uses open-source datasets, such as Spider, MultiWOZ, Huatuo-26M, and ShenNong-TCM. The MultiWOZ dataset is widely used in research on task-oriented dialogue. The Huatuo-26M dataset is collected from publicly accessible data without personal information and is available to academic researchers. The ShenNong-TCM dataset is generated by GPT-3.5 based on a traditional Chinese medicine knowledge graph. As a result, these datasets should not be regarded as controversial. All interactions are generated by LLMs, which may inevitably include hallucinations or incorrect information. Human evaluators are also fully aware that they are reading interactions generated by LLMs. We use LLMs to assist with paper writing by handling language-level tasks such as grammar checking and revision.
Reproducibility Statement
@@ Appendix A Model list
Appendix B Converage analysis
The training curve of prompt optimisation based on different settings (e.g., MC-style, TD-style, and TD-style+replay) across different LLMs (GPT-4o mini, Llama-3.1-8B, Llama-3.1-70B, and Llama-4-scout) is shown in Figure 6 (The result of Gemini-2.0-flash is shown in 4(a) previously).
The training curves become stable after epoch 3 (trained with 300 dialogues), and the TD-style+replay setting improves the stability. However, since existing LLMs are not batch-invariant, which means their behaviour will be impacted by different batch sizes, there is unavoidable variance caused by their nondeterministic behaviour (He & Thinking Machines Lab, 2025).
@@ Appendix B Converage analysis
Appendix C Prompts
@@
C.1 An example of the system prompt before and after optimisation by RPO
Figure 11 shows the original prompt of FnCTOD and Figure 12 is the prompt optimised by $\text{RPO}_{\text{TD+replay}}$.
Figure 11: The system prompt of FnCTOD before prompt optimisation.

### v2

@@ 5.2 Effect of different styles and input signals of textual-based feedbacker
The training curves of FnCTOD optimised by the methods of MC-style, TD-style, and TD-style+replay with Gemini-2.0-flash are shown in Figure 4 (See results with other LLMs in Figure 8). Similar to the behaviour in traditional RL optimisation, MC-style exhibits higher variance during the early stages of training, whereas TD-style is more stable and converges faster. With further training, their final performances become comparable. In contrast, incorporating experience replay into the rewriter yields more stable training and achieves the best overall performance. A qualitative analysis is presented in Appendix D, including the difference between the TD- and MC-style feedbacks and the corresponding optimised system prompts.
We conduct a further ablation study on the impact of different information as input to the feedbacker (as shown in Figure 5). The basic setting passes the dialogue in pure text. The subjective setting includes the user goal, and the believe setting adds the API call in comparison to the basic setting, respectively. The full setting is our proposed TD-style+replay, including both the user goal and the system API call.
Figure 5: The training curves of different info for RPO.
@@
Figure 6: The impact of noisy external verifiable evaluation signal.
@@
To substantiate the impact of the quality of textual feedback more, we flipped $20\%$, $40\%$, and $60\%$ of the external evaluation signal to study the robustness to noisy external evaluation signals. The experiment is conducted with Gemini-2.0-flash on task-oriented dialogue, and the result is as shown in Figure 6.
The performance is robust when the noise level is up to $40\%$. The optimisation ability of RPO declines when the noise reaches $60\%$, which reverses the reward signal on average and naturally harms any optimisation. On the other hand, if the noise level is less than $50\%$, its performance is close to the clean setting. In other words, the turn-level feedback, which is not conditioned on the external feedback, provides a learning signal that, to some extent, mitigates the noise in the external evaluation signal.
5.3 Prompting limitations on underrepresented topics in LLMs
In general medicine, our method consistently outperforms the fully fine-tuned HuatuoGPT-II with an 86.7% win rate and is preferred over other prompting-based baselines (see 7(a)). On the other hand, traditional Chinese medicine is more challenging. For example, our system’s preference rate drops by 41% compared to Huatuo when transitioning from general medicine to traditional Chinese medicine (see 7(b)). However, despite this drop in preference, our proposed method is still favoured in general.
@@ 6 Conclusions
Impact Statement
This paper presents work aimed at advancing the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be highlighted here. To be more specific, this work uses open-source datasets, such as Spider, MultiWOZ, Huatuo-26M, and ShenNong-TCM. The MultiWOZ dataset is widely used in research on task-oriented dialogue. The Huatuo-26M dataset is collected from publicly accessible data without personal information and is available to academic researchers. The ShenNong-TCM dataset is generated by GPT-3.5 based on a traditional Chinese medicine knowledge graph. As a result, these datasets should not be regarded as controversial. All interactions are generated by LLMs, which may inevitably include hallucinations or incorrect information. Human evaluators are also fully aware that they are reading interactions generated by LLMs. We use LLMs to assist with paper writing by handling language-level tasks such as grammar checking and revision.
Software and Data
@@ Appendix A Model list
Appendix B RPO algorithm
The pseudo-code of RPO is given in Algorithm 1.
Algorithm 1 Reinforced Prompt Optimisation (RPO)
1:Input: initial prompt $\textit{prompt}^{1}$; training data $\mathcal{D}_{\text{train}}$; validation data $\mathcal{D}_{\text{val}}$;
2:Input: system $\textup{LLM}_{\text{sys}}$; user $\textup{LLM}_{\text{usr}}$;
3:Input: external evaluator $\texttt{Eval}()$; simulation environment $\texttt{Interact}()$;
4:Input: feedbacker $\textup{LLM}_{F}$; rewriter $\textup{LLM}_{R}$;
5:Input: number of epochs NN; trajectories $n_{\text{train}},n_{\text{val}}$; rewritten prompt candidates $n_{p}$.
6:for $i=1$ to NN do
7:
8:  Phase 1: Trajectory Collection
9:  for $j=1$ to $n_{\text{train}}$ do
10:   $t\leftarrow\texttt{Interact}(\textup{LLM}_{\text{sys}}(\cdot|\textit{prompt}^{i}),\textup{LLM}_{\text{usr}}(\cdot|\text{goal}\sim\mathcal{D}_{\text{train}}))$
11:   $\mathcal{T}\leftarrow\mathcal{T}\cup\{t\}$
12:  end for
13:
14:  Phase 2: TD-style Feedback Generation
15:  for all $t\in\mathcal{T}$ do
16:   for $t_{j}$ in trajectory $t$ do
17:    $\textit{feedback}_{\textup{TD},j}^{i}\leftarrow\textup{LLM}_{F}(t_{1},\textit{feedback}_{\textup{TD},1}^{i},\dots,t_{j})$ {turn-level feedback, Eqn. 2}
18:   end for
19:   $\text{s}\leftarrow\texttt{Eval}(t)$
20:   $\textit{feedback}_{\textup{TD}(t)}^{i}\leftarrow\textup{LLM}_{F}(\textit{feedback}_{\textup{TD},1}^{i},\textit{feedback}_{\textup{TD},2}^{i},\ldots,\text{s})$ {dialogue-level feedback}
21:  end for
22:  $\textit{feedback}_{\textup{TD}}^{i}\leftarrow\textup{LLM}_{F}(\textit{feedback}_{\textup{TD}(\mathcal{T})}^{i})$ {final feedback}
23:
24:  Phase 3: Prompt Rewriting and Evaluation
25:  for $k=1$ to $n_{p}$ do
26:   Generate new prompts: $\textit{prompt}^{i+1,k}\leftarrow\textup{LLM}_{R}(\textit{prompt}^{i},\textit{feedback}_{\textup{TD}}^{i},\cdots,\textit{prompt}^{1},\textit{feedback}_{\textup{TD}}^{1})$ {Eqn. 5}
27:
28:   for $j=1$ to $n_{\text{val}}$ do
29:    $t\leftarrow\texttt{Interact}(\textup{LLM}_{\text{sys}}(\cdot|\textit{prompt}^{i+1,k}),\textup{LLM}_{\text{usr}}(\cdot|\text{goal}\sim\mathcal{D}_{\text{val}}))$
30:    $\mathcal{T}^{k}_{\text{val}}\leftarrow\mathcal{T}^{k}_{\text{val}}\cup\{t\}$
31:   end for
32:   $\text{success}^{k}\leftarrow\texttt{Eval}(\mathcal{T}^{k}_{\text{val}})$ {evaluate candidate prompt}
33:  end for
34:
35:  Phase 4: Policy Update
36:  $k^{*}\leftarrow\arg\max_{k}\text{success}^{k}$
37:  $\textit{prompt}^{i+1}\leftarrow\textit{prompt}^{i+1,k^{*}}$
38:end for
39:Return $\textit{prompt}^{N+1}$
Appendix C Converage analysis
The training curve of prompt optimisation based on different settings (e.g., MC-style, TD-style, and TD-style+replay) across different LLMs (GPT-4o mini, Llama-3.1-8B, Llama-3.1-70B, and Llama-4-scout) is shown in Figure 8 (The result of Gemini-2.0-flash is shown in Figure 4 previously).
The training curves become stable after epoch 3 (trained with 300 dialogues), and the TD-style+replay setting improves the stability. However, since existing LLMs are not batch-invariant, which means their behaviour will be impacted by different batch sizes, there is unavoidable variance caused by their nondeterministic behaviour (He and Thinking Machines Lab, 2025).
@@ Appendix C Converage analysis
Appendix D Qualitative Analysis of TD-style feedback
One example of the turn-level feedback generation by $\text{RPO}_{\textup{TD}+\text{replay}}$ is shown in Figure 9. After receiving the first turn $t_{1}$, $\textit{feedback}_{TD,1}$ is generated, including a positive user emotion estimation and a prediction that the conversation will be successful since there is no mistake at $t_{1}$. However, the system makes a mistake in turn $t_{2}$, where the API call includes wrong information, i.e. the user mentions “Tuesday”, but the system puts “sunday” in the API call (highlighted in red), resulting in a not found result. $\textit{feedback}_{\textup{TD},2}$ points it out, includes negative user emotion prediction, and suggests the system should acknowledge the user’s request properly (highlighted in yellow). This result demonstrates that our proposed feedbacker can estimate the user satisfaction and user success, and provide suggestions properly with reasoning.
The turn-level feedbacks will be summarised by the TD-style feedbacker into a dialogue-level feedback (as shown in Figure 10 and referred to Line 20 in Algorithm 1), and then all dialogue-level feedbacks from the training set will be summarised as the final feedback for this epoch (as shown in Figure 11 and referred to Line 22 in Algorithm 1).
D.1 Comparison with the MC-style feedback
The feedback produced by both the TD-style feedbacker (Figure 11) and the MC-style feedbacker (Figure 12) can offer useful guidance, such as preventing conversational loops or handling cases where information cannot be found. However, the TD-style feedbacker captures more fine-grained signals and provides more detailed suggestions, which in turn leads to a faster convergence, as mentioned in Section 5.2.
Figure 9: The turn-level feedback generated by the TD-style feedbacker. The first blue block is the instruction of the feedbacker (see full prompt in Figure 19). The mistake made by the system is highlighted in red, and the suggestion by the feedbacker is highlighted in yellow, manually.
Figure 10: The turn-level feedback in Figure 9 will be summarised by the TD-style feedbacker into a dialogue-level feedback.
Figure 11: All the dialogue-level feedback, e.g. Figure 10, will be further summarised by the TD-style feedbacker into the final feedback $\textit{feedback}_{\textup{TD}}$ of this epoch. The suggestions related to goal tracking, API calls, and prevention of looping are manually highlighted in yellow, red, and green, respectively.
Figure 12: The feedback generated by the MC-style feedbacker.
D.2 The system prompt before and after optimisation
Figure 13 shows the original prompt of FnCTOD, and Figure 14 and Figure 15 are the prompt optimised by MC-style and $\text{RPO}_{\text{TD+replay}}$, respectively.
Based on the optimisation results, the system prompt optimised by $\text{RPO}_{\text{TD+replay}}$ includes more details to deal with multi-turn task-oriented dialogue, including how to deal with domain switching.
Figure 13: The system prompt of FnCTOD before prompt optimisation.
Figure 14: The system prompt of FnCTOD after it is optimised by MC-style for 8 epochs. MC-style is built with Gemini-2.0-Flash. The format is generated by the rewriter in markdown format. For illustration, the instructions of goal tracking (yellow) and looping prevention (green) are manually highlighted.
Appendix E Prompts